the length of a rectangle is 4cm more than the width and the perimeter is at least 48 cm . what are the smallest possible dimensions for the rectangle
Answer by graphmatics(170)

Let w be the width then w+4 is the length and 2*w+2*(w+4)=4*w+8 is the expression for the perimeter. So
4*w+8>=48
4*w>=40
w>=10